could someone please help me solve this equation and solve by the substitution method?
The solution to the system of equations below is (8,4)
3x - 5y = 4
7x + 3y = 68
thank you

Answer by Alan3354(69443)

3x - 5y = 4
7x + 3y = 68
---------------
9x - 15y = 12 Eqn 1 times 3
35x+ 15y = 340 Eqn 2 times 5 --> 15y = 340 - 35x
Sub for 15y in Eqn 1 times 3
-----------
9x - (340-35x) = 12
44x = 352
x = 8
-----------
3*8 - 5y = 4
y = 4
----------
If you solve for x or y as a 1st step, you get into messy fractions.
Make it easy on yourself.